OpenStudy (nincompoop):

proof that √2 is irrational assuming that √2 = p/q, where p and q are integers not both even

OpenStudy (nincompoop):

winding up with 2q^2 = p^2 makes p^2 an even number

OpenStudy (zzr0ck3r):

ok so 2=(p/q)^2 so 2=p^2/q^2 2q^2=p^2 so we know p is even because p^2 is even now 2q^2 = (2k)^2 = 4k^2 q^2 = 2k^2 so q is even this is a contradiction to (p,q) = 1 thus sqrt(2) is irrational

OpenStudy (zzr0ck3r):

assume q^2 is even and q is odd q^2 = (2k+1)^2 = 4k^2+4k+1 = 2(2k^2+2k)+1 = odd a contradction to q^2 is even so q^2 is even imples q is even

OpenStudy (zzr0ck3r):

now please don't murder me for this

OpenStudy (anonymous):

My favorite way of proving this statement is using the Rational Roots theorem.

OpenStudy (anonymous):

Second favorite way is using the Fundamental Theorem of Arithmetic.

OpenStudy (nincompoop):

that's what I was trying to do next @joemath314159 but it seems tedious

OpenStudy (anonymous):

Thats not very precise, can i see your proof anyway out of boredem

OpenStudy (anonymous):

Which one, the Rational roots? or Fundamental Theorem? I think they are both shorter than the standard proof by contradiction.

OpenStudy (anonymous):

the fundamental theorem 1

OpenStudy (anonymous):

Sure. So the Fundamental theorem of Arithmetic is the one that says we can factor any integer uniquely into primes. So I'm going to prove \(\sqrt{2}\) is irrational by contradiction with that theorem. Assume \(\sqrt{2}=\frac{a}{b}\) for some integers a and b, \(q\ne 0\). Then \(2b^2=a^2\). By the Fundamental Theorem of Arithmetic, we can factor a as \(a=p_1p_2\cdots p_n\) and b as \(b=q_1q_2\cdots q_m.\) Therefore:\[2b^2=a^2\Longrightarrow 2(q_1q_2\cdots q_m)^2=(p_1p_2\cdots p_n)^2.\]Now we count the number of primes on each side. There are \(2m+1\) primes of the left hand side (the +1 because of the 2 hanging out in the front), and \(2n\) on the right. Since m and n are natural numbers, an odd number can never equal an even. This is the contradiction.

OpenStudy (anonymous):

er, that shouldnt be \(q\ne 0\), it should be \(b\ne 0\). my bad >.<

OpenStudy (nincompoop):

it's the same ending as @zzr0ck3r indicated

OpenStudy (anonymous):

The parity of m and n themselves isn't relevant. Its the parity of 2m+1 and 2n.

OpenStudy (anonymous):

Oh I see in oder for them to be equal they must have the same number of prime factors which would require that $$2m=2n+1$$

OpenStudy (anonymous):

Exactly :)

OpenStudy (nincompoop):

p and q shouldn't have a factor in common?

OpenStudy (anonymous):

yes

OpenStudy (nincompoop):

the contradiction winds up both p and q having he same factor of 2

OpenStudy (anonymous):

Thats correct :)

OpenStudy (nincompoop):

I think I know what you did there, @joemath314159 but shouldn't you have assumed that p and q also have to be in lowest terms?

OpenStudy (zzr0ck3r):

@Jack199 what is not percise?

OpenStudy (anonymous):

With the F. T. A. approach, you aren't aiming to contradict anything dealing with reduced fraction \(\frac{p}{q}\). So it doesn't matter if the fraction is in lowest terms or not.

OpenStudy (anonymous):

I wasn't responding to anything you wrote I was responding to joe, but then he clarified with a proof.

OpenStudy (zzr0ck3r):

ahh. I think this proof (i think this is how one of the Pythagoreans proved it) is the best proof because it relies on nothing but knowledge of parity, and relatively primness.

OpenStudy (anonymous):

Its a classic for sure.

OpenStudy (nincompoop):

I remember my proof in 4th grade went along the lines that any number in fractional form, if the quotient would seem to end up as non-terminating, it would fit the description of irrational if you attempted to finish it. LUL the person is the one who isn't rational ;)

OpenStudy (anonymous):

The textbook standard for learning how to prove things by contradiction.

OpenStudy (nincompoop):

thanks both @joemath314159 and @zzr0ck3r

OpenStudy (zzr0ck3r):

I think this and the proof that there are infinite primes, using the idea of dividing all the primes + 1 with a prime, are my favorite proofs

OpenStudy (anonymous):

Indeed. Its so elegant.

OpenStudy (anonymous):

I like things that are proved with the pigeonhole principle.

OpenStudy (nincompoop):

square root of any prime number is irrational?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (anonymous):

Yes. In fact, the prove i presented using the F.T.A only works if we are dealing with the square root of the prime. The method zzr0ck3r presented works for any case.

OpenStudy (anonymous):

proof*

OpenStudy (zzr0ck3r):

if a*a = prime then a|p

OpenStudy (nincompoop):

what the heck

OpenStudy (zzr0ck3r):

there are probably 20+ ways to prove it.

OpenStudy (zzr0ck3r):

there are over 100 documented proofs for the pythagorean theorem One of them was done by President Garfield and its actually one of the cooler ones.

OpenStudy (nincompoop):

mind boggled

OpenStudy (zzr0ck3r):

the infinite primes one is like this assume there are finite primes {p,p_1,p_2....} P = the product of them all + 1 P = p_1*p_2....p_n+1 so P is not prime because its the product of all the primes + 1 so P is divisible by some prime = p because P is composite so p|P but p cant divide our product of primes because then it would be in the product, so it must divide 1. A contradiction I love this.

OpenStudy (nincompoop):

I appreciate the time both of you have poured into this seemingly "popular" and "classic" proof problem.